OpenStudy (anonymous):

What is the perimeter of a rectangle with a length of 8 and a width of 2?

OpenStudy (anonymous):

\[8+8+2+2\]

OpenStudy (anonymous):

/[8+8+2+2/]

OpenStudy (anonymous):

Perimeter = 2 x length + 2 x width. Perimeter = 2(8) + 2(2) = 16 + 4 = 20